OpenStudy (anonymous):

The average cost per slice of pizza is 88 cents with a standard deviation of 5 cents. What is the z-score of a slice of pizza with a cost of 99 cents? 0.45 1.16 2.2 7.4

OpenStudy (whpalmer4):

z-score is the number of standard deviations away from the average. how far is 99 cents away from 88 cents? what do you get when you divide that number by the standard deviation?

OpenStudy (anonymous):

99-88/5=2.2?

OpenStudy (anonymous):

so C

OpenStudy (anonymous):

yes that is correct

OpenStudy (anonymous):

thank you! i need more help

OpenStudy (anonymous):

Which type of bias is best described by the statement, “the way in which the survey is written favors a particular response”? Response Non-Response Wording Freedom

OpenStudy (anonymous):

i was thinking A?

OpenStudy (anonymous):

@Lethal

OpenStudy (anonymous):

@fiberdust

OpenStudy (anonymous):

@ybarrap

OpenStudy (whpalmer4):

wording bias is where the wording of the question biases the response.

OpenStudy (ybarrap):

I would say Wording

OpenStudy (anonymous):

What is the probability of choosing a 2 or a face card from a deck of cards? 0.057 0.077 0.25 0.3077

OpenStudy (anonymous):

i looked over the last question and yes it is wording A ^

OpenStudy (anonymous):

there are 4 cards in a deck with a 2 and 3 face cards per suit(4suits) = 12 so 12/52 + 4/52 = 16/52 would be .307692308

OpenStudy (anonymous):

@fiberdust need help one more

OpenStudy (anonymous):

Consider the lengths of alligators to be { 18.7, 12.3, 18.6, 16.4, 15.7, 18.3, 15.8, 14.9, 17.6, 16.4, 16.7, 17.8, 16.2, 13.8, 15.2, 14.7, 13.2, 15.8, 14.6, 16.6, 16.2, 18.1, 17.5, 17.3 }. According to the Empirical Rule, 99.7% of the data should fall between what two numbers? 11.1 and 21.3 12.3 and 18.1 12.8 and 19.6 13.4 and 18.8

OpenStudy (anonymous):

not familiar with this type of problem but my guess would be the number 12.3 is smallest and choosing 18.1 as ur leaving out 18.7 and 18.3

OpenStudy (anonymous):

i dont have a calculator for this one but im guessing its c? Find the standard deviation for the following data: { 20 , 30 , 38 , 40 } 3.416 4.570 7.133 9.092

OpenStudy (whpalmer4):

empirical rule sales 99.7% falls within 3 standard deviations of the mean.

OpenStudy (anonymous):

@whpalmer4 we were agreeing that the alligator question would be b?

OpenStudy (anonymous):

not c 7.133 for std deviation if u take ur mean /4 u get 32 then take each number for example (20 -32)^2+ .... (40-32)^2 then u should get 248 then ur n-1 would be 3 so 248/3 and sqrt it u get 9.092

OpenStudy (whpalmer4):

The mean of that data set is 16.2. The standard deviation is 1.7.

OpenStudy (whpalmer4):

16.2 +/- 3*1.7

OpenStudy (anonymous):

Thank you! @fiberdust @whpalmer4 you guys are amazing!

OpenStudy (anonymous):

your welcome hope u understand it better